OpenStudy (anonymous):

please ASAP distance between a to b is 120 miles.car travels from a to b @ 60 miles per hr and returns from b to a along the same route @40 miles per hr.what is the average speed of the round trip?

OpenStudy (abhisar):

\(\huge\text {Average speed}\) = \(\huge\frac{Total~distance}{Total~time~taken}\)

OpenStudy (abhisar):

Time taken in going from a to b = 120/60=2hr

OpenStudy (abhisar):

Time taken in coming back = 120/40=3hr

OpenStudy (abhisar):

Total time taken in complete round trip= 2+3=5hr, Total distance travelled in complete round trip= 120*2=240

OpenStudy (abhisar):

Average speed=240/5=48miles/hr

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

answer is 56 as given that made me confused this is what i too presumed

OpenStudy (abhisar):

Ir should be 48km/hr

OpenStudy (abhisar):

*it

OpenStudy (anonymous):

i too think so lets c the result tomoro hope it works thanx bro

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (abhisar):

Indian ?

OpenStudy (anonymous):

nop NEPALI

OpenStudy (anonymous):

neighbours

OpenStudy (abhisar):

Which book ?

OpenStudy (anonymous):

not a book model test question fo a fren i had to solve it

OpenStudy (abhisar):

ok